# Don’t Pay Attention

### v1

@@ Appendix B Activation or No Activation in the Enricher
In this study, we illustrate the performance of Avey with and without an activation function in the enricher. To this end, we trained a model with 145 million parameters using 10 billion tokens from the Fineweb dataset [41]. The model employs an expansion factor of 4×4$\times$4 × in the enricher (i.e., each embedding dimension is expanded fourfold by the enricher), a tail size of 50% (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm [108] as a normalization technique, and no activation function in the contextualizer. Additionally, the context width (i.e., the maximum number of tokens that can be input to and processed by the contexutalizer simultaneously) is set to 1024 and a constant learning rate of $1e{-3}$ is maintained throughout training.
The study excludes the ranker and focuses solely on the neural processor. Besides, it evaluates four activation functions, namely, GELU [33], ReLU [62], ReLU2 [10], and SiLU [76]. All other experimental settings follow the methodology detailed in Appendix A. Table 3 summarizes the results. As shown, ReLU2 yielded an improvement in performance versus a baseline with no activation, hence, was adopted as the default activation function for the enricher throughout our experiments presented in Sections 3.3 and 3.4. It is important to note, however, that the lowest perplexity was provided by GELU and not ReLU2 (although the difference in perplexity was very minimal). While perplexity quantifies how effectively the model predicts the next token in the training dataset, it remains a proxy for overall modeling capability and does not always precisely predict downstream task performance.
Appendix C Activation or No Activation in the Contextualizer
Table 4: Avey’s performance without and with an activation function in the contextualizer. The study involves only the neural processor within Avey, trains it on 10B tokens, and uses ReLU2 within the enricher, capitalizing on the results shown in Table 3.
@@ Appendix C Activation or No Activation in the Contextualizer
We now evaluate Avey with and without an activation function in the contextualizer. We use the same experimental settings outlined in Appendix B and add to that ReLU2 as an activation function in the enricher, capitalizing on the findings therein. We also experiment with four activation functions, namely, GELU, ReLU, ReLU2, and SiLU. Results are summarized in Table 4. As illustrated, the best performance was achieved without any activation function in the contextualizer, thus was employed as the default configuration in all our experiments reported in Sections 3.3 and 3.4.
Appendix D What is the Best Expansion Factor?
Table 5: The effect of the expansion factor on Avey’s performance. The study involves only the neural processor within Avey, trains it on 10B tokens, adopts ReLU2 within the enricher, and does not use an activation function within the contextualizer, building upon the results portrayed in Tables 3 and 4.
@@ Appendix D What is the Best Expansion Factor?
In this study, we vary the expansion factor in the enricher, defined as the degree to which each input embedding is expanded. Specifically, we evaluate several expansion factors, ranging from $2x$ to $8x$ as shown in Table 5, while keeping the total model parameter count constant (e.g., with a $2x$ expansion factor we use 34 layers, while with a $4x$ one we utilize 20 layers). The experimental setup follows the methodology outlined in Appendix B, but incorporates ReLU2 as an activation function in the enricher and omits any activation function in the contextualizer, aligning with the findings reported in Appendices B and C. As depicted in the table, an expansion factor of $4x$ yielded the best performance, hence, was set as the default configuration in the enricher throughout all our experiments presented in Sections 3.3 and 3.4.
Appendix E What is the Best Tail Size?
Table 6: The effect of the tail size on Avey’s performance. The study involves only the neural processor within Avey, trains it on 10B tokens, adopts ReLU2 within the enricher, does not use an activation function within the contextualizer, and utilizes an expansion factor of $4x$, as recommended by the findings demonstrated in Tables 34, and 5.
@@ Appendix E What is the Best Tail Size?
We now examine the impact of forwarding a tail portion of each enriched embedding to the contextualizer. More precisely, we vary the size of this tail portion, referred to as the tail size, from 10% to 90% of each enriched embedding, as illustrated in Table 6. The study follows the experimental setup described in Appendix B and employs ReLU2 as an activation function in the enricher, no activation function in the contextualizer, and an expansion factor of $4x$, based on the findings presented in Appendices B, C, and D, respectively. As depicted in Table 6, the best performance was accomplished using a tail size of 50%, thus was adopted as the default configuration for Avey in all our experiments reported in Sections 3.3 and 3.4.
Appendix F Deeper Models and Narrower Embeddings, or Shallower Models and Wider Embeddings
Table 7: Avey’s performance across different model configurations, including wider embedding dimensions (e.g., 1536 under 0.5B-parameter model) with shallower layers (e.g., 24 layers under 0.5B-parameter model), or narrower embedding dimensions (e.g., 768 under 0.5B-parameter model) with deeper layers (e.g., 90 layers under 0.5B-paramter model). The models with 140M, 0.5B, and 1.5B parameters are referred to as small Avey, medium Avey, and large Avey in the text.
@@ Appendix F Deeper Models and Narrower Embeddings, or Shallower Models and Wider Embeddings
Table 8: Avey’s performance under different sequence lengths, $N$, split sizes, $S$, and top-$k$ values. All models (a total of 69) were trained on 10B tokens using 140 million parameters. The sweet spot in terms of downstream task performance was at $N=512$, $S=64$, and $k=7$; hence, it was adopted as Avey’s default configuration.
$N$ $S$ $k$ Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
The objective of this study is to determine whether a narrower embedding dimension with a greater model depth yields better or worse performance than a wider embedding dimension with fewer layers. The study utilizes the experimental setup described in Appendix B and leverages the findings presented in Appendices BCD, and E. Consequently, it utilizes ReLU2 as an activation function in the enricher, no activation function in the contextualizer, an expansion factor of $4x$, and a tail size of 50%.
To begin with, we evaluated a small Avey model (referred to as small Avey) with 140 million parameters, using three different embedding dimensions, 512, 768, and 1024. These configurations resulted in 40, 20, and 11 layers, respectively, to maintain a constant parameter count. Table 7 summarizes the results. As illustrated, the configuration with an embedding dimension of 768 and a layer count of 20 outperformed the other two configurations.
Afterwards, we assessed a larger Avey model with 500 million parameters (referred to as medium Avey), using three different embedding dimensions, 768, 1024, and 1536. These configurations resulted in 90, 54, and 24 layers, respectively, while keeping the total parameter count constant. As shown in Table 7, the setup with an embedding dimension of 768 and 90 layers yielded the best performance among the three tested ones.
Finally, we examined an even larger Avey model with 1.5 billion parameters (referred to as large Avey), using three different embedding dimensions, 1536, 2048, and 2560. To maintain a constant parameter count across configurations, these dimensions corresponded to 80, 48, and 30 layers, respectively. As portrayed in Table 7, the configuration with an embedding dimension of 2048 and 48 layers delivered the best performance among the three considered configurations.
The above results suggest a trend, whereby wider embedding dimensions (e.g., 1024 in small Avey; 1563 in medium Avey; and 2560 in large Avey) paired with shallower architectures (e.g., 11 layers in small Avey; 24 layers in medium Avey; and 30 layers in large Avey) tend to underperform deeper models (e.g., 40 and 20 layers in small Avey; 90 and 54 layers in medium Avey; and 80 and 48 layers in large Avey) with narrower embeddings (e.g., 512 and 768 in small Avey; 768 and 1024 in medium Avey; and 1536 and 2048 in large Avey). As such, in all our experiments discussed in Sections 3.3 and 3.4, we employed deeper models with narrower embedding dimensions, namely, the best performing small Avey, medium Avey, and large Avey in Table 7.
Interestingly, Table 7 also highlights that certain benchmarks benefit more from increased model capacity than others. For instance, a commonsense reasoning benchmark like HellaSwag demonstrates performance improvements of 20.36% and 28.5% under 0.5B-parameter and 1.5B-parameter models, respectively, compared to a 140M-parameter baseline. In contrast, a question-answering benchmark such as SIQA exhibits only a modest gain of 2.4% under both 0.5B-parameter and 1.5B-parameter models relative to the 140M-parameter baseline, suggesting less sensitivity to model size.
Appendix G What are the Best Sequence Length, Split Size, and Top-$k$ Values?
@@
We now analyze how Avey’s perplexity and overall performance are affected by variations in three key parameters, the ranker’s top-$k$ selected splits, the split size $S$, and the sequence length $N$. We consider three values for $N$, $256$, $512$, and $1024$. For each $N$, the split size $S$ is grown geometrically, starting from $16$ and doubling at each step, up to the maximum permissible value $N/2$. Subsequently, for any $N$ and $S$, the number of top-ranked splits $k$ can range from $1$ (i.e., contextualizing the current split with one additional relevant split) up to $N/S-1$. To tame the quantity of experiments, we increase $k$ arithmetically from $1$ to a maximum of $15$, whenever possible, using an increment of $2$. Finally, we use the experimental configurations recommended in Appendices BCDE, and F.
@@
As shown in Table 8, Avey’s perplexity is highest when both $S$ and $k$ are very small (e.g., $S=16$ and $k=1$). While a small $S$ can help filter out irrelevant embeddings and denoise contextualization, pairing it with a very small $k$ can deprive the contextualizer of sufficient context to build expressive representations [^9] Recall from Section 2.2.2 that the context width $C$ is defined as $C=S(k+1)$.. To expand the context (i.e., increase its width $S(k+1)$) and enrich the resulting embeddings, either $S$ or $k$ can be increased. For example, as we increased $k$ under $N=256$ and $S=16$, perplexity decreased and benchmark performance improved. However, a larger context does not always yield better outcomes, especially when involving a high proportion of irrelevant embeddings. This behavior was observed when $k$ was increased beyond $3$ under $N=1024$ and $S=128$, resulting in higher perplexity and diminished downstream accuracy.
@@
In contrast to $S$ and $k$, the sequence length $N$ determines the size of the candidate pool from which the ranker selects the top-$k$ splits for any current split during training. A larger $N$ allows the ranker to reach farther back in the sequence history, potentially retrieving more relevant splits and lowering perplexity. For example, increasing $N$ from $512$ to $1024$, while holding $S=16$ and $k=15$ constant, reduced perplexity by $5.4\%$. Nevertheless, lower perplexity does not always translate into improved downstream performance. For instance, with $N=512$ and $S=16$, $k=15$ yielded the lowest perplexity, yet $k=5$ achieved higher benchmark accuracy. As noted in Appendix B, the loss remains only a proxy for overall modeling capability and does not always exactly predict downstream task performance. As demonstrated in Table 8, the best empirical performance was obtained with $N=512$, $S=64$, and $k=7$, which was, accordingly, adopted as Avey’s default configuration.
Appendix H RMSNorm or LayerNorm
@@
Table 9: Avey’s performance with RMSNorm and LayerNorm. A model with 153 million parameters was trained on 10B tokens using both the neural processor and ranker with the best configuration from Table 8.
@@
In all previous runs across the appendices, we used RMSNorm [108] as a normalization technique. We now test Avey with another standard normalization method, namely, LayerNorm [1]. To begin with, we note that Avey normalizes input embeddings before each layer (as illustrated in Fig. 3) and output embeddings once after the final layer and prior to token prediction.
We evaluate each normalization technique using Avey’s complete architecture, including its neural processor and ranker. The ranker is configured according to the best-performing setting identified in Appendix G (i.e., sequence length $N=512$, split size $S=64$, and number of top-ranked splits $k=7$). The neural processor employs the optimal configurations reported in Appendices BCDE, and F. As before, we trained the resulting model with 153 million parameters on 10 billion tokens from the Fineweb dataset. Table 9 summarizes the results. As shown, RMSNorm slightly outperforms LayerNorm on average, and is therefore adopted in Avey and used consistently across all the experiments presented in Sections 3.3 and 3.4.
Appendix I Learning Rate: To Decay or Not to Decay?
Table 10: Avey’s performance with two types of schedules, constant learning rate (LR) and cosine decay, starting from different peak learning rates. All models involved the neural processor and ranker, and were trained with 153 million parameters on 10B tokens, using both the neural processor and ranker with the best configuration from Table 8.
@@ Appendix I Learning Rate: To Decay or Not to Decay?
In all previous runs across the appendices, we used a constant learning rate of $1e{-3}$. In this study, we evaluate the performance of Avey under varying maximum learning rates and learning rate schedules. Specifically, we compare two schedules, cosine decay and constant learning rate (which can be viewed as cosine decay with an infinite cycle length, effectively eliminating any decay). The study adopts the experimental configurations suggested in Appendices BCDE, and F. In addition, it employs Avey’s complete architecture, including its neural processor and ranker, with the best configuration from Table 8.
@@
As shown in Table 10, cosine decay consistently achieves lower losses than the constant schedule across the tested learning rates. This observation aligns with findings from the Chinchilla paper [35], which indicates that when the cosine cycle length significantly exceeds the total number of training steps (by at least 25%), model performance tends to deteriorate. Notably, the longest cycle length arises when the schedule is constant. In contrast, setting it to approximately match the training duration yields the best final loss [35].

### v2

@@ 3.2.1 Activation or No Activation in the Enricher
In this study, we evaluate Avey’s performance with and without an activation function in the enricher. We train a 145M-parameter model on 10B tokens from the FineWeb dataset (fineweb2023). The configuration uses an enricher expansion factor of $4\times$ (each embedding is expanded fourfold), a tail size of $50\%$ (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm (zhang2019rmsnorm) for normalization, and no activation in the contextualizer. The context width (i.e., the maximum number of tokens processed concurrently by the contextualizer) is set to 1024, and we employ a constant learning rate of 1e31e{-3} throughout training.
This study excludes the ranker and focuses solely on the neural processor. Besides, it evaluates four activation functions in the enricher, namely, GELU (hendrycks2016gelu), ReLU (nair2010relu), $\mathrm{ReLU}^{2}$ (chowdhery2022palm), and SiLU (ramachandran2017swish). All other experimental settings follow the methodology in Section 3.1. Results are summarized in Table 3. As shown, $\mathrm{ReLU}^{2}$ yields a clear accuracy improvement over the no-activation baseline and is therefore adopted as the default enricher activation for Avey. Notably, GELU attains the lowest perplexity, though the gap relative to $\mathrm{ReLU}^{2}$ is minimal. While perplexity measures next-token predictive quality on the training distribution, it is an indirect proxy for modeling capability and does not always precisely predict downstream task performance.
3.2.2 Activation or No Activation in the Contextualizer
Table 4: Avey’s performance without and with an activation function in the contextualizer. The study isolates the neural processor (without the ranker), trains on 10B tokens, and fixes $\mathrm{ReLU}^{2}$ in the enricher, following the findings in Table 3.
@@ 3.2.2 Activation or No Activation in the Contextualizer
Next, we evaluate Avey both with and without an activation function in the contextualizer. We employ the same experimental setup as in Section 3.2.1 and, consistent with the findings reported therein, fix $\mathrm{ReLU}^{2}$ as the activation for the enricher. We then assess four candidate activations for the contextualizer, namely, GELU, ReLU, $\mathrm{ReLU}^{2}$, and SiLU. The results, summarized in Table 4, indicate that the best performance is achieved without any activation in the contextualizer. Accordingly, we adopt this configuration as Avey’s default setting.
3.2.3 What is the Best Expansion Factor?
Table 5: The effect of the expansion factor on Avey’s performance. The study isolates the neural processor (without the ranker), trains on 10B tokens, fixes $\mathrm{ReLU}^{2}$ as the enricher activation, and uses no activation in the contextualizer, following the configurations established in Tables 3 and 4.
@@ 3.2.3 What is the Best Expansion Factor?
In this study, we investigate the effect of the expansion factor in the enricher, defined as the degree to which each input embedding is expanded. Specifically, we evaluate expansion factors ranging from $2\times$ to $8\times$ (see Table 5) while maintaining a fixed total parameter budget (e.g., $2\times$ uses 34 layers, whereas $4\times$ uses 20 layers). The experimental setup follows that of Section 3.2.1, employing $\mathrm{ReLU}^{2}$ in the enricher (chowdhery2022palm) and no activation in the contextualizer, consistent with the findings from Sections 3.2.1 and 3.2.2. As shown in Table 5, an expansion factor of $4\times$ delivers the strongest performance and is therefore adopted as the default configuration for the enricher.
3.2.4 What is the Best Tail Size?
Table 6: The effect of the tail size on Avey’s performance. The study isolates the neural processor (without the ranker), trains on 10B tokens, fixes $\mathrm{ReLU}^{2}$ as the enricher activation, uses no activation in the contextualizer, and adopts an expansion factor of $4\times$, following the findings in Tables 3, 4, and 5.
@@ 3.2.4 What is the Best Tail Size?
We now examine the impact of forwarding a tail portion of each enriched embedding to the contextualizer. Concretely, we vary the tail size (or the fraction of each enriched embedding passed to the contextualizer) from $10\%$ to $90\%$ (see Table 6). The study follows the setup outlined in Section 3.2.1, using $\mathrm{ReLU}^{2}$ in the enricher, no activation in the contextualizer, and an expansion factor of $4\times$, consistent with the findings in Sections 3.2.1, 3.2.2, and 3.2.3. As depicted in Table 6, a tail size of $50\%$ yields the strongest performance and is accordingly employed as Avey’s default configuration.
3.2.5 Deeper Models and Narrower Embeddings, or Shallower Models and Wider Embeddings
Table 7: Avey’s performance across configurations that trade embedding width for depth, including wider embeddings with fewer layers (e.g., 1536-dimensional embeddings with 24 layers for the 0.5B-parameter model) versus narrower embeddings with more layers (e.g., 768-dimensional embeddings with 90 layers for the 0.5B-parameter model). Throughout the text, the 140M-, 0.5B-, and 1.5B-parameter models are referred to as small Avey, medium Avey, and large Avey, respectively.
@@ 3.2.5 Deeper Models and Narrower Embeddings, or Shallower Models and Wider Embeddings
The objective of this study is to determine whether a narrower embedding dimension with a greater model depth yields better or worse performance than a wider embedding dimension with fewer layers. The study utilizes the experimental setup described in Section 3.2.1 and leverages the findings of Sections 3.2.13.2.23.2.3, and 3.2.4. Consequently, it employs $\mathrm{ReLU}^{2}$ as the activation function in the enricher, no activation in the contextualizer, an expansion factor of $4\times$, and a tail size of $50\%$.
We begin by evaluating a 140M-parameter Avey model (referred to as small Avey) with three embedding dimensions, 512, 768, and 1024. To maintain a constant parameter budget, these dimensions yielded 40, 20, and 11 layers, respectively. Results are summarized in Table 7. As shown, the configuration with a 768-dimensional embedding and 20 layers outperforms the other two.
We next evaluate a 500M-parameter Avey model (denoted as medium Avey) with embedding dimensions of 768, 1024, and 1536. To keep the total parameter count fixed, these settings resulted in 90, 54, and 24 layers, respectively. As reported in Table 7, the configuration with a 768-dimensional embedding and 90 layers achieves the strongest performance among the three.
Lastly, we examine a 1.5B-parameter Avey model (referred to as large Avey) with embedding dimensions of 1536, 2048, and 2560. Again, to maintain a constant parameter budget, these settings yielded 80, 48, and 30 layers, respectively. As illustrated in Table 7, the configuration with a 2048-dimensional embedding and 48 layers delivers the best performance among the three.
The above results suggest a trend, whereby wider embedding dimensions (e.g., 1024 in small Avey; 1563 in medium Avey; and 2560 in large Avey) paired with shallower architectures (e.g., 11 layers in small Avey; 24 layers in medium Avey; and 30 layers in large Avey) tend to underperform deeper models (e.g., 40 and 20 layers in small Avey; 90 and 54 layers in medium Avey; and 80 and 48 layers in large Avey) with narrower embeddings (e.g., 512 and 768 in small Avey; 768 and 1024 in medium Avey; and 1536 and 2048 in large Avey). As such, in all our subsequent experiments, we adopt deeper models with narrower embedding dimensions, namely, the best performing small Avey, medium Avey, and large Avey in Table 7.
Interestingly, Table 7 indicates that some benchmarks benefit more from increased model capacity than others. For example, the commonsense reasoning task HellaSwag shows gains of $20.36\%$ and $28.5\%$ under the 0.5B- and 1.5B-parameter models, respectively, relative to the 140M-parameter baseline. In contrast, the question–answering benchmark SIQA exhibits only a modest improvement of $2.4\%$ for both the 0.5B- and 1.5B-parameter models, suggesting lower sensitivity to model size.
3.2.6 What are the Best Sequence Length, Split Size, and Top-$k$ Values?
Table 8: Avey’s performance under varying sequence lengths (NN), split sizes ($S$), and top-$k$ values. The study trains 69 models with 140M parameters on 10B tokens. The strongest downstream performance is achieved at $N=512$, $S=64$, and $k=7$.
NN $S$ $k$ Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@ 3.2.6 What are the Best Sequence Length, Split Size, and Top-$k$ Values?
We now analyze how Avey’s perplexity and overall performance vary with three key parameters, the ranker’s top-$k$ selected splits, the split size $S$, and the sequence length $N$. We consider $N\in\{256,512,1024\}$. For each $N$, the split size $S$ grows geometrically from $16$, doubling at each step up to the maximum permissible value $N/2$. Subsequently, for any $(N,S)$, the number of top-ranked splits $k$ ranges from $1$ (i.e., contextualizing the current split with one additional relevant split) up to N/S1N/S-1. To keep the experiment count tractable, we vary $k$ arithmetically from $1$ to at most $15$, in increments of $2$ whenever feasible. All runs use the best configurations established in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5.
@@
As shown in Table 8, Avey’s perplexity is highest when both $S$ and $k$ are very small (e.g., $S=16$ and $k=1$). Although a small $S$ can help filter out irrelevant embeddings and denoise contextualization, pairing it with a very small $k$ can starve the contextualizer of sufficient signal to form expressive representations [^11] Recall from Section 2.2.2 that the context width $C$ is defined as $C=S(k+1)$.. To expand the effective context (i.e., increase $S(k+1)$) and enrich the resulting representations, one can increase either $S$ or $k$. For example, under $N=256$ and $S=16$, increasing $k$ monotonically reduces perplexity and improves benchmark accuracy. However, a larger context does not invariably yield better outcomes, especially when it admits a higher proportion of irrelevant embeddings. This effect appears when $k>3$ for $N=1024$ and $S=128$, where perplexity rises and downstream accuracy degrades.
@@
Unlike $S$ and $k$, the sequence length $N$ controls the size of the candidate pool from which the ranker selects top-$k$ splits for the current split during training. A larger $N$ allows the ranker to reach further back in the sequence history, potentially retrieving more relevant splits and lowering perplexity. For example, increasing $N$ from $512$ to $1024$ while holding $S=16$ and $k=15$ fixed reduces perplexity by $5.4\%$. Nevertheless, lower perplexity does not always translate into higher downstream accuracy. With $N=512$ and $S=16$, for instance, $k=15$ attains the lowest perplexity, yet $k=5$ yields better benchmark accuracy. As noted in Section 3.2.1, loss remains a proxy for overall modeling capability and does not always precisely predict downstream performance. As shown in Table 8, the strongest empirical results are obtained with $N=512$, $S=64$, and $k=7$, which we accordingly adopt as Avey’s default configuration.
3.2.7 RMSNorm or LayerNorm
@@
Table 9: Avey’s performance with RMSNorm versus LayerNorm. The study trains a 153M-parameter model on 10B tokens using the full architecture (neural processor + ranker) with the best configuration from Table 8.
@@
In all preceding studies, we employed RMSNorm (zhang2019rmsnorm) as the normalization method. We now compare Avey against an alternative standard, that is, LayerNorm (ba2016layer). Recall that Avey applies pre-layer normalization to the input embeddings at every layer (as illustrated in Fig. 3) and a single post-layer normalization to the output embeddings after the final layer, immediately before token prediction.
We evaluate each normalization method using Avey’s full architecture, comprising both the neural processor and the ranker. The ranker is fixed to the best-performing setting from Section 3.2.6 (i.e., sequence length $N=512$, split size $S=64$, and top-$k=7$). The neural processor adopts the optimal configurations identified in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5. As before, we train a 153M-parameter model on 10B tokens from the FineWeb dataset. Results are summarized in Table 9. On average, RMSNorm slightly outperforms LayerNorm and is subsequently used in Avey by default.
3.2.8 Learning Rate: To Decay or Not to Decay?
Table 10: Avey’s performance under two learning-rate schedules, constant and cosine decay, across varying peak learning rates. All models use the full architecture (neural processor + ranker) with the best configuration from Table 8, and are trained with 153M parameters on 10B tokens.
@@ 3.2.8 Learning Rate: To Decay or Not to Decay?
In all prior experiments, we used a constant learning rate of 1e31e{-3}. In this study, we assess Avey’s performance under varying peak learning rates and learning-rate schedules. Specifically, we compare a cosine-decay schedule against a constant schedule (i.e., cosine decay with effectively infinite cycle length, yielding no decay). The study adopts the experimental configurations established in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5. In addition, we utilize Avey’s full architecture, encompassing the neural processor and ranker, with the best configuration from Table 8.
@@
As shown in Table 10, cosine decay consistently achieves lower loss than a constant schedule across the learning rates tested. This trend aligns with the Chinchilla findings (hoffmann2022training), which report that when the cosine cycle length substantially exceeds the total number of training steps (by at least $25\%$), performance degrades. Notably, a constant schedule corresponds to an effectively infinite cycle length. In contrast, setting the cycle length to approximately match the training duration yields the best final loss (hoffmann2022training).